Ministry of Higher Education (MOHE) put a guideline that maximum of 30% students are allowed to be in campus to control the spread of Covid-19. In average, 13 students will come to Statistics class with standard deviation of 3 students. A sample of 20 students are chosen. Find probability that the average number of students come to Statistics class is less than 15?
1
Expert's answer
2020-10-22T18:13:16-0400

"P(\bar X<15)=P(Z<\frac{15-13}{\frac{3}{\sqrt{20}}})=P(Z<2.98)=0.9986."